# Can LLMs Beat Classical Hyperparameter Optimization Algorithms? A Study on autoresearch

### 29 Mar 2026 v2

@@ Abstract
The autoresearch repository enables an LLM agent to search for optimal hyperparameter configurations on an unconstrained search space by editing the training code directly. Given a fixed compute budget and constraints, we use autoresearch as a testbed to compare classical hyperparameter optimization (HPO) algorithms against LLM-based methods on tuning the hyperparameters of a small language model. Within a fixed hyperparameter search space, classical HPO methods such as CMA-ES and TPE consistently outperform LLM-based agents. However, an LLM agent that directly edits training source code in an unconstrained search space narrows the gap to classical methods substantially despite using only a self-hosted open-weight 27B model. Methods that avoid out-of-memory failures outperform those with higher search diversity, suggesting that reliability matters more than exploration breadth. While small and mid-sized LLMs struggle to track optimization state across trials, classical methods lack domain knowledge. To bridge this gap, we introduce Centaur, a hybrid that shares CMA-ES’s internal state, including mean vector, step-size, and covariance matrix, with an LLM111Named after the mythological half-human, half-horse hybrid: our method merges LLM reasoning with classical optimization.. Centaur achieves the best result in our experiments, with its 0.8B variant outperforming the 27B variant, suggesting that a cheap LLM suffices when paired with a strong classical optimizer. The 0.8B model is insufficient for unconstrained code editing but sufficient for hybrid optimization, while scaling to 27B provides no advantage for fixed search space methods. Preliminary experiments with the frontier model Gemini 3.1 Pro Preview do not close the gap to classical methods. Code is available at https://github.com/ferreirafabio/autoresearch-automl.
@@ 1 Introduction
autoresearch (Karpathy, 2025a) demonstrated that an LLM agent can iteratively edit training code to improve a small language model (50M parameters), reaching a val_bpb of 0.978 in less than 100 trials on a single H100. Building on that, Ziv (2025) showed that TPE, a classical AutoML method, with expert-designed hyperparameters can outperform Karpathy’s agent within a similar budget. Hyperparameter optimization (HPO) is a central component of AutoML (Feurer and Hutter, 2019; Bischl et al., 2023). These results raise two questions: (i) how do other classical HPO methods perform on this task? and (ii) can LLM-based HPO methods outperform classical ones?
To answer these questions, we benchmark 9 HPO methods, spanning 4 classical, 4 LLM-based, and 1 hybrid, on Karpathy’s autoresearch task. All methods operate under the same 24-hour GPU training budget with 3 seeds. To reduce human priors, we automatically extract 14 hyperparameters from the training script; while the ranges require some domain knowledge, the HP selection itself is automated, removing manual search space curation. Figure˜1 compares all 9 methods using 27B LLM variants against cumulative training wall-time, showing that classical methods find better configurations than LLM-based agents within the fixed search space. The exception is Karpathy Agent (Code), which directly edits training source code and is competitive with classical methods, even though the latter find a similar performing hyperparameter configuration faster. All LLM methods use a self-hosted open-weight model (Qwen3.5-27B); preliminary experiments with the frontier model Gemini 3.1 Pro Preview (Section˜A.2) show competitive but not superior performance.
@@
We introduce Centaur, a hybrid that shares CMA-ES’s full internal state with the LLM and achieves the best result in our experiments.
We analyze search diversity, OOM rates, and model scaling across all methods. Preliminary experiments with the frontier model Gemini 3.1 Pro Preview suggest that simply scaling the LLM optimizer does not close the gap to classical methods.
@@ 2 Related Work
Classical HPO spans a wide range of approaches (Bischl et al., 2023; Feurer and Hutter, 2019), from random search (Bergstra and Bengio, 2012) and Bayesian optimization with Gaussian process surrogates (Snoek et al., 2012) to sequential model-based approaches with random forests such as SMAC (Hutter et al., 2011), tree-structured Parzen estimators such as TPE (Bergstra et al., 2011), and evolution strategies such as CMA-ES (Hansen, 2016). We focus on single-task HPO methods without multi-fidelity or transfer to isolate each optimizer’s ability to learn the landscape from scratch. Methods such as Hyperband (Li et al., 2017), BOHB (Falkner et al., 2018), transfer HPO (Wistuba and Grabocka, 2021; Arango et al., 2024), and zero-shot approaches (Öztürk et al., 2022) could in principle be applied to this benchmark but are not in the scope of this study.
@@ 3 Experimental Setup
Our main experiments use Qwen3.5 (Team, 2026) (0.8B and 27B) as the LLM optimizer, self-hosted via vLLM (Kwon et al., 2023) on the same GPU that trains the optimizee (the 50M-parameter language model). We additionally run frontier model experiments with Gemini 2.5 Flash (Comanici et al., 2025), 3.1 Flash-Lite, and 3.1 Pro Preview (Google DeepMind, 2026) via the Gemini API. To ensure equal resource allocation, we capped training VRAM to 80 GB for all methods, comparable to the H100 used by Karpathy (Karpathy, 2025a), and reserved the remaining memory for the vLLM server when using self-hosted models. We disabled Qwen3.5’s thinking mode and sampled with temperature 0.7, limiting outputs to 2048 tokens for fixed-HP methods and 16384 tokens for code editing. All prompt templates are available in our repository.
@@
| LLAMBO (Optuna) | Fixed | OptunaHub port: binary surrogate labels, random categorical sampling |
@@ 4 Centaur: CMA-ES Guided LLM Optimization
Centaur shares CMA-ES’s full internal state with the LLM on a fraction of trials. On every trial, CMA-ES proposes a candidate configuration from its multivariate Gaussian, parameterized by mean vector , step-size , and covariance matrix . On 30% of trials, the LLM receives CMA-ES’s proposal along with , , , the top-5 configurations, and the last 20 trials.
The LLM may override the proposal, and in practice does so in nearly all cases: 100% of the time with the 27B model and 95% with the 0.8B model. Crucially, CMA-ES updates its internal state from all trial results, including those where the LLM overrode its proposal, so the optimizer continuously learns from the full trajectory.
@@ 5.1 Classical methods outperform LLMs in fixed search spaces
We report both LLAMBO variants because the OptunaHub implementation differs from the original paper in how it handles surrogate labels, categorical HPs, and failed trials, which substantially affects OOM rates and performance (see Section˜A.8 for details).
@@ 5.2 Unconstrained code editing is viable but requires model scale
Scaling the LLM from 0.8B to 27B is essential for unconstrained code editing but provides no advantage for fixed-HP optimization, as can be seen in Figure˜2: 0.8B is insufficient for unconstrained code editing (Karpathy Agent Code: 0.9910 vs 0.9814 with 27B), while fixed-HP methods saw no benefit from scaling (Karpathy Agent 14 HPs: 0.9904 vs 0.9908). All LLM methods in Figures˜1 and 2 used open-weight Qwen3.5. Preliminary comparisons with Gemini 2.5 Flash (Comanici et al., 2025) and Gemini 3.1 Flash-Lite (Google DeepMind, 2026) as the LLM optimizer (Section˜A.2) show no improvement over Qwen3.5-27B. Gemini 3.1 Flash-Lite fails to generate valid code in 87–94% of trials for unconstrained editing, but performs comparably to Qwen3.5-0.8B in the fixed search space, confirming that code editing is where model capability matters most. Gemini 3.1 Pro Preview (Figure˜3) is competitive with Qwen3.5-27B for both unconstrained code editing and Centaur, with lower OOM rates for code editing (5% vs 12%), but does not close the gap to the best classical methods. Gemini 3.1 Pro Preview results are based on a single seed; with the frontier models tested so far, we do not see evidence that simply scaling the LLM optimizer shifts the balance in favor of LLM-based approaches.
@@ 5.3 Hybrid optimization: best of both worlds
Centaur outperformed all methods including CMA-ES alone by using the LLM on only 30% of trials. As described in Section˜4, the LLM receives CMA-ES’s full internal state (, , ), the top-5 configurations, and the last 20 trials, and almost always overrides CMA-ES’s proposal (100% for 27B, 95% for 0.8B). Despite constituting only 30% of trials, LLM trials contributed 25% of incumbent improvements (Section˜A.5), confirming that the LLM provides complementary value beyond what CMA-ES finds alone. Beyond improving the mean, Centaur substantially reduces CMA-ES’s cross-seed variance: std drops from 0.0036 for CMA-ES alone to 0.0005 for Centaur, suggesting that the LLM stabilizes the optimizer by injecting domain knowledge that prevents unfavorable seeds from drifting. Notably, Centaur [0.8B] outperformed Centaur [27B], suggesting that a cheap LLM suffices when paired with a strong classical optimizer. This stands in contrast to unconstrained code editing, where the 0.8B model is insufficient: when the classical optimizer handles the search trajectory, the LLM’s role appears to reduce to refining promising candidates, a task that does not require the same level of model capability as generating code edits from scratch. We ablate the LLM ratio (see Appendix Tables˜4 and 8). For the 0.8B model, achieves the best result (0.9735), slightly outperforming the default (0.9776), while degrades to 0.9789. For the 27B model, performs best (0.9745), but collapses to 0.9903, which is worse than CMA-ES alone. This confirms that CMA-ES should retain majority control of the optimization trajectory: the LLM is most effective as an occasional informed perturbation, not as the primary search driver.
@@ 6 Conclusion
We benchmarked classical, LLM-based, and hybrid HPO methods on small-scale language model training under the same budgets and constraints. Within the fixed, constrained search space, classical HPO methods such as CMA-ES and TPE consistently outperformed LLM-based agents, with OOM avoidance emerging as a stronger predictor of performance than search diversity. Restricting LLMs to a fixed hyperparameter search space does not leverage their strengths; the one setting where a pure LLM method competes with classical approaches is unconstrained code editing (Karpathy, 2025a), where the agent directly modifies training source code and narrows the gap to classical methods substantially despite using only a self-hosted open-weight model. However, this requires sufficient model scale: the 0.8B model is insufficient for unconstrained code editing, while it suffices for hybrid optimization. Centaur achieved the best result in our experiments by using the LLM on only 30% of trials, preserving fast convergence from CMA-ES while benefiting from LLM-informed suggestions.
Our study evaluated a single task with open-weight models (Qwen3.5 0.8B and 27B). Preliminary single-seed experiments with frontier model Gemini 3.1 Pro Preview shows competitive but not clearly superior performance to Qwen3.5-27B (Section˜A.2). More experiments and other models are needed to determine whether frontier models can close the remaining gap to classical methods. Also, all LLM methods in this benchmark operate without tools or external knowledge retrieval. Equipping LLM optimizers with agentic capabilities such as paper search, documentation lookup, or code analysis tools is a promising direction. Additionally, exploring other classical optimizers as the hybrid base and pairing CMA-ES with a code-editing LLM agent could allow the search space to co-evolve with the optimization trajectory.
@@ A.1 Convergence by Trial Number
A.2 Frontier Model Comparison
To test whether a stronger LLM optimizer changes the balance between classical and LLM-based methods, we ran Karpathy Agent (Code) and Centaur with three Gemini variants: Gemini 2.5 Flash [Comanici et al., 2025], Gemini 3.1 Flash-Lite [Google DeepMind, 2026], and Gemini 3.1 Pro Preview [Google DeepMind, 2026]. Gemini 2.5 Flash and 3.1 Flash-Lite runs were stopped early after 16–18 hours of cumulative training time as no meaningful improvement was observed beyond that point. Gemini 3.1 Pro Preview was run for the full 24-hour budget but limited to a single seed due to API cost constraints.
Figures˜6 and 7 compare the Flash variants with Qwen3.5-27B, but neither outperforms it. Notably, Karpathy Agent (Code) with Gemini 3.1 Flash-Lite suffers from extremely high failure rates (87–94%), as the model frequently generates code modifications that cause the training script to crash (93% of failures are runtime errors, not OOM). In the fixed search space, however, Gemini 3.1 Flash-Lite performs comparably to Qwen3.5-0.8B. Figure˜3 shows that Gemini 3.1 Pro Preview is competitive with Qwen3.5-27B for both unconstrained code editing (Karpathy Agent Code: 0.9827 vs 0.9814, 5% OOM) and Centaur in the fixed search space. However, these are single-seed results; Gemini Pro initially shows more exploratory behavior with a higher early OOM rate, proposing deeper models (DEPTH15), before converging to performance comparable to Qwen3.5-27B.
A.3 Centaur LLM Ratio Ablation

### 13 Apr 2026 v4

@@ Abstract
The autoresearch repository enables an LLM agent to optimize hyperparameters by editing training code directly. We use it as a testbed to compare classical HPO algorithms against LLM-based methods on tuning the hyperparameters of a small language model under a fixed compute budget. When defining a fixed search space over autoresearch, classical methods such as CMA-ES and TPE consistently outperform LLM-based agents, where avoiding out-of-memory failures matters more than search diversity. Allowing the LLM to directly edit source code narrows the gap to the classical methods but does not close it, even with frontier models available at the time of writing such as Claude Opus 4.6 and Gemini 3.1 Pro Preview. We observe that LLMs struggle to track optimization state across trials. In contrast, classical methods lack the domain knowledge of LLMs. To combine the strengths of both, we introduce Centaur, a hybrid that shares CMA-ES’s interpretable internal state, including mean vector, step-size, and covariance matrix, with an LLM111Named after the mythological half-human, half-horse hybrid: our method merges LLM reasoning with classical optimization.. Centaur achieves the best result in our experiments, and a 0.8B LLM already suffices to outperform all classical and pure LLM methods. Unconstrained code editing requires larger models to be competitive with classical methods. We further analyze search diversity, model scaling from 0.8B to frontier models, and ablate the fraction of LLM-proposed trials in Centaur. All in all, our results suggest that LLMs are most effective as a complement to classical optimizers, not as a replacement. Code is available at https://github.com/ferreirafabio/autoresearch-automl & interactive demo at https://ferreirafabio.github.io/autoresearch-automl.
@@ 1 Introduction
autoresearch (Karpathy, 2025a) demonstrated that an LLM agent can iteratively edit training code to improve a small language model (50M parameters), reaching a val_bpb of 0.978 in less than 100 trials on a single H100. HPO is a central component of AutoML (Feurer and Hutter, 2019; Bischl et al., 2023) that can encompass various hyperparameters of a learning pipeline, from model architecture and optimizer settings to data augmentation strategies (Arango et al., 2024; Wagner et al., 2022). This raises two questions: (i) how do other classical HPO methods perform on this task? and (ii) can LLM-based HPO methods outperform classical ones?
To answer these questions, we benchmark 9 HPO methods, spanning 4 classical, 4 LLM-based, and 1 hybrid, on Karpathy’s autoresearch task. All methods operate under the same 24-hour GPU training budget with 3 seeds. To reduce human priors, we automatically extract 14 hyperparameters from the training script; while the ranges require some domain knowledge, the HP selection itself is automated, removing manual search space curation. Figure˜1 compares all 9 methods using 27B LLM variants against cumulative training wall-time, showing that classical methods find better configurations than LLM-based agents within the fixed search space. The exception is Karpathy Agent (Code), which directly edits training source code and is competitive with classical methods, even though the latter find a similar performing hyperparameter configuration faster. All LLM methods use a self-hosted open-weight model (Qwen3.5-27B); experiments with frontier models Gemini 3.1 Pro Preview (Google DeepMind, 2026) and Claude Opus 4.6 (Anthropic, 2026) show that pure LLM agents do not outperform classical methods even with frontier models, though scaling benefits hybrid methods more than pure LLM agents (Figure˜2).
@@
We introduce Centaur, a hybrid that shares CMA-ES’s full internal state with the LLM and achieves the best result in our experiments, where a 0.8B LLM already suffices to outperform all classical and pure LLM methods.
We analyze search diversity, OOM rates, and model scaling across all methods. Experiments with frontier models confirm that simply scaling the LLM does not outperform classical methods, though Centaur with Opus 4.6 extends its lead further.
@@ 2 Related Work
Classical HPO spans a wide range of approaches (Bischl et al., 2023; Feurer and Hutter, 2019), from random search (Bergstra and Bengio, 2012) and Bayesian optimization with Gaussian process surrogates (Snoek et al., 2012) to sequential model-based approaches with random forests such as SMAC (Hutter et al., 2011), tree-structured Parzen estimators such as TPE (Bergstra et al., 2011), and evolution strategies such as CMA-ES (Hansen, 2016). We focus on single-task HPO methods without multi-fidelity or transfer to isolate each optimizer’s ability to learn the landscape from scratch. Methods such as Hyperband (Li et al., 2017), BOHB (Falkner et al., 2018), transfer HPO (Wistuba and Grabocka, 2021; Arango et al., 2024), zero-shot approaches (Öztürk et al., 2022), multi-task transfer (Strangmann et al., 2024), and hyperparameter-robust training (Ferreira et al., 2022) could in principle be applied to this benchmark but are not in the scope of this study.
@@ 3 Experimental Setup
Our main experiments use Qwen3.5 (Team, 2026) (0.8B and 27B) as the LLM optimizer, self-hosted via vLLM (Kwon et al., 2023) on the same GPU that trains the optimizee (the 50M-parameter language model). We additionally run frontier model experiments with Gemini 2.5 Flash (Comanici et al., 2025), 3.1 Flash-Lite, and 3.1 Pro Preview (Google DeepMind, 2026) via the Gemini API. To ensure equal resource allocation, we capped training VRAM to 80 GB for all methods, comparable to the H100 used by Karpathy (Karpathy, 2025a), and reserved the remaining memory for the vLLM server when using self-hosted models. We disabled Qwen3.5’s thinking mode and sampled with temperature 0.7, limiting outputs to 2048 tokens for fixed-HP methods and 16384 tokens for code editing. Each LLM-based method receives a prompt containing the optimization goal (minimize val_bpb), the model class (GPT-2 scale transformer with Muon+AdamW optimizer), the dataset (FineWeb), the hardware (H200 GPU with VRAM budget and OOM warning), the search space with bounds, and the trial history including both successful and failed configurations. Full prompt templates for all five LLM-based methods are reproduced verbatim in Section˜A.11.
@@
| LLAMBO (Optuna) | Fixed | OptunaHub (Shibata et al., 2025) port: binary surrogate labels, random categorical sampling |
@@ 4 Centaur: CMA-ES Guided LLM Optimization
Centaur shares CMA-ES’s full internal state with the LLM on a fraction of trials. On every trial, CMA-ES proposes a candidate configuration from its multivariate Gaussian, parameterized by mean vector , step-size , and covariance matrix . On 30% of trials, the LLM receives CMA-ES’s proposal along with , , , the top-5 configurations, and the last 20 trials. The LLM may override the proposal, and in practice does so in nearly all cases: 100% of the time with the 27B model and 95% with the 0.8B model. Crucially, CMA-ES updates its internal state from all trial results, including those where the LLM overrode its proposal, so the optimizer continuously learns from the full trajectory.
@@ 5.1 Classical methods outperform LLMs in fixed search spaces
We report both LLAMBO variants because the OptunaHub implementation differs from the original paper in how it handles surrogate labels, categorical HPs, and failed trials, which substantially affects OOM rates and performance (see Section˜A.9 for details).
@@ 5.2 Unconstrained code editing is viable but requires model scale
Scaling the LLM does not consistently close the gap. Comparisons with Gemini 2.5 Flash (Comanici et al., 2025) and Gemini 3.1 Flash-Lite (Google DeepMind, 2026) as the LLM optimizer (Section˜A.3) show no improvement over Qwen3.5-27B. Gemini 3.1 Flash-Lite fails to generate valid code in 87–94% of trials for unconstrained editing, but performs comparably to Qwen3.5-0.8B in the fixed search space, confirming that code editing is where model capability matters most. Gemini 3.1 Pro Preview (Figure˜2) is competitive with Qwen3.5-27B for both unconstrained code editing (0.9826 vs 0.9814) and Centaur (0.9767 vs 0.9763), with lower OOM rates for code editing (2% vs 12%), but does not close the gap to the best classical methods. Claude Opus 4.6 (Anthropic, 2026) narrows the gap for code editing (0.9773 vs 0.9814 with Qwen3.5-27B), making it competitive with CMA-ES (0.9785), but still falls short of TPE (0.9768). For pure LLM code editing, scaling the model helps but does not surpass the best classical methods.
Scaling the LLM from 0.8B to 27B is essential for unconstrained code editing but provides no advantage for fixed-HP optimization, as can be seen in Figure˜3: 0.8B is insufficient for unconstrained code editing (Karpathy Agent Code: 0.9910 vs 0.9814 with 27B), while fixed-HP methods saw no benefit from scaling (Karpathy Agent 14 HPs: 0.9904 vs 0.9908). All LLM methods in Figures˜1 and 3 used open-weight Qwen3.5.
@@ 5.3 Hybrid optimization: best of both worlds
Centaur outperformed all methods including CMA-ES alone by using the LLM on only 30% of trials. As described in Section˜4, the LLM receives CMA-ES’s full internal state (, , ), the top-5 configurations, and the last 20 trials, and almost always overrides CMA-ES’s proposal (100% for 27B, 95% for 0.8B). Despite constituting only 30% of trials, LLM trials contributed 25% of incumbent improvements (Section˜A.6), confirming that the LLM provides complementary value beyond what CMA-ES finds alone. Beyond improving the mean, Centaur substantially reduces CMA-ES’s cross-seed variance: std drops from 0.0036 for CMA-ES alone to 0.0005 for Centaur, suggesting that the LLM stabilizes the optimizer by injecting domain knowledge that prevents unfavorable seeds from drifting. Within the open-weight models tested, Centaur [0.8B] outperformed Centaur [27B], and Gemini 3.1 Pro Preview performed on par with both (0.9766, 0.9763, 0.9767), suggesting a capability plateau where a cheap LLM suffices when paired with a strong classical optimizer. However, Claude Opus 4.6 (Anthropic, 2026) breaks through this plateau: Centaur [Opus 4.6] achieves 0.97390.0012, improving over the open-weight variants by 0.0024 and establishing a new benchmark leader (Figure˜2). This improvement is not explained by better OOM avoidance (17% vs 15% for Qwen 27B) or more trials, but by higher-quality suggestions during the LLM’s 30% of turns. This indicates that the plateau is not fundamental: a sufficiently capable frontier model provides additional gains even in the hybrid setting, though the practical benefit of open-weight models remains that they achieve competitive performance at a fraction of the inference cost. We ablate the LLM ratio (see Appendix Tables˜4 and 11). For the 0.8B model, achieves the best result (0.9735), slightly outperforming the default (0.9776), while degrades to 0.9789. For the 27B model, performs best (0.9745), but collapses to 0.9903, which is worse than CMA-ES alone. This confirms that CMA-ES should retain majority control of the optimization trajectory: the LLM is most effective as an occasional informed perturbation, not as the primary search driver.
@@ 6 Conclusion
We benchmarked classical, LLM-based, and hybrid HPO methods on small-scale language model training under identical budgets. Within a fixed search space, classical methods consistently outperformed LLM-based agents, with OOM avoidance mattering more than search diversity. Restricting LLMs to a fixed hyperparameter search space does not leverage their strengths; allowing the LLM to directly edit source code narrows the gap but does not close it, even with frontier models such as Claude Opus 4.6 and Gemini 3.1 Pro Preview. Centaur achieved the best result in our experiments by using the LLM on only 30% of trials, where a 0.8B LLM already suffices for hybrid optimization while unconstrained code editing requires larger models. Centaur with Opus 4.6 extends this lead further, though the improvement comes from higher-quality suggestions rather than better OOM avoidance. All in all, our results suggest that LLMs are most effective as a complement to classical optimizers, not as a replacement.
Our study evaluated a single task with open-weight models (Qwen3.5 0.8B and 27B) and two frontier models (Gemini 3.1 Pro Preview and Claude Opus 4.6). Scaling from 0.8B to 27B to Gemini Pro yields no improvement for hybrid optimization, but Centaur with Opus 4.6 breaks through this plateau. More benchmarking and experiments are needed to determine the generality of these findings. All LLM methods in this benchmark operate without tools or external knowledge retrieval. Equipping LLM optimizers with agentic capabilities such as paper search, documentation lookup, or code analysis tools is a promising direction. Additionally, exploring other classical optimizers as the hybrid base and pairing CMA-ES with a code-editing LLM agent could allow the search space to co-evolve with the optimization trajectory.
@@ A.1 Convergence by Trial Number
A.2 Per-Seed Convergence
Figures˜6, 7 and 8 show the same convergence plots as the main text but with individual seed trajectories (thin, transparent) overlaid on the mean (thick, solid) instead of std bands. This reveals cross-seed variance and whether the mean is driven by one outlier seed or consistent behavior.
A.3 Frontier Model Comparison
To test whether a stronger LLM optimizer changes the balance between classical and LLM-based methods, we ran Karpathy Agent (Code) and Centaur with three Gemini variants: Gemini 2.5 Flash [Comanici et al., 2025], Gemini 3.1 Flash-Lite [Google DeepMind, 2026], and Gemini 3.1 Pro Preview [Google DeepMind, 2026]. Gemini 2.5 Flash and 3.1 Flash-Lite runs were stopped early after 16–18 hours of cumulative training time as no meaningful improvement was observed beyond that point. Gemini 3.1 Pro Preview was run for the full 24-hour budget with 3 seeds.
Figures˜9 and 10 compare the Flash variants with Qwen3.5-27B, but neither outperforms it. Notably, Karpathy Agent (Code) with Gemini 3.1 Flash-Lite suffers from extremely high failure rates (87–94%), as the model frequently generates code modifications that cause the training script to crash (93% of failures are runtime errors, not OOM). In the fixed search space, however, Gemini 3.1 Flash-Lite performs comparably to Qwen3.5-0.8B. Figure˜2 shows that Gemini 3.1 Pro Preview is competitive with Qwen3.5-27B for both unconstrained code editing (Karpathy Agent Code: 0.98260.0005 vs 0.98140.0057, 2% OOM) and Centaur in the fixed search space (0.97670.0016 vs 0.97630.0006). Gemini Pro achieves lower variance than Qwen for code editing but does not improve the mean, confirming that scaling the LLM optimizer does not close the gap to classical methods.
A.4 Centaur LLM Ratio Ablation